OpenStudy (anonymous):

14.)15-g=23-2g 15.)5y+1.8=4y-3.2 16.)6a-5=4a+2 17.)4y-8-2y+5=0 18.)6(n-4)=3n 19.)5(2-g)=0 I REALLY HAVE NOT A CLUE?????? alg 2

OpenStudy (anonymous):

please help??? im struggling???

thomaster (thomaster):

What do you need to do? solve them?

OpenStudy (anonymous):

yupp. im struggling w/it

OpenStudy (anonymous):

if our helping start at 15 plzzz

thomaster (thomaster):

Oke \(\large 5y+1.8=4y-3.2\) What is the first thing to do

OpenStudy (anonymous):

subtract 1.8?

OpenStudy (anonymous):

??????

thomaster (thomaster):

yes subtract 1.8 on both sides. \(\large 5y\cancel{+1.8}=4y-3.2-1.8\) \(\large 5y=4y-5\)

thomaster (thomaster):

So what's the next step?

OpenStudy (anonymous):

divide 4y?

thomaster (thomaster):

No, subtract 4y on both sides \(\large 5y-4y=\cancel{4y}−5\) what do we have now?

OpenStudy (anonymous):

1y=-5 you divide by 1 y=-5

thomaster (thomaster):

Very good :) So that's the answer. Now the next one: \(\large 6a-5=4a+2\) what's the first step?

OpenStudy (anonymous):

subtract 4a??

thomaster (thomaster):

You could do that, but usually we start with bringing the numbers to the right. Then the letters to the left. When you'd start with the letter (a in this case) you need to subtract 4a as you said. \(\large 6a-4a-5=\cancel{4a}+2\) \(\large 2a-5=2\) You probably know what to do next

OpenStudy (anonymous):

o wait no I didn't. I did mine like yours, my bad

thomaster (thomaster):

So what's the next thing you do? :P

OpenStudy (anonymous):

+5?????????? next

thomaster (thomaster):

yes It'll be \(\large 2a=7\) now only 1 thing left

OpenStudy (anonymous):

a=3 1/2

OpenStudy (anonymous):

divide by 2.. a=3and1/2

thomaster (thomaster):

Yep, so \(\large a = 3\frac{1}{2}\) or 3.5

thomaster (thomaster):

Could you do \(\large 4y-8-2y+5=0\) ?

OpenStudy (anonymous):

mind walking me thru a few more pleaseee???? I haven't passed this alg 1 test and im a senior in alg 2 I HAve to know this stuff before they get harder, pleaseee help w/more???

thomaster (thomaster):

What is the first thing to do? \(\large 4y−8−2y+5=0\)

OpenStudy (anonymous):

would the first line be: 2y-3=0????

thomaster (thomaster):

exactly, you can combine like terms first :) Now what?

OpenStudy (anonymous):

idk?

thomaster (thomaster):

we have \(2y-3=0\) now, You need to bring the numbers to the right side of the equation.

OpenStudy (anonymous):

so...???

thomaster (thomaster):

So you add 3 on both sides \(2y\cancel{-3}=0+3\) \(2y=3\)

OpenStudy (anonymous):

1and1/2 answer?

thomaster (thomaster):

yes \(\dfrac{3}2\to1\dfrac{1}{2}\)

thomaster (thomaster):

So y=1.5

OpenStudy (anonymous):

help with 18-26????? pleaseeeeeeee

thomaster (thomaster):

18-26??

OpenStudy (anonymous):

distribute 6

OpenStudy (anonymous):

18 through 26

thomaster (thomaster):

oh oke, first 18: \(\large 6(n-4)=3n\)

thomaster (thomaster):

and yes distribute 6 :)

OpenStudy (anonymous):

3 word story problems. 4 equations

OpenStudy (anonymous):

subtract 3n?

thomaster (thomaster):

No first distribute 6

thomaster (thomaster):

What do you get when you distribute 6?

OpenStudy (anonymous):

I did that. 6n-24=3n then subtract 3n next???

thomaster (thomaster):

Yes, what will the equation be when you subtract 3n on both sides?

OpenStudy (anonymous):

6n-3n-24=3n-3n

OpenStudy (anonymous):

3n-24=0?

thomaster (thomaster):

exactly :) now bring the number to the right

OpenStudy (anonymous):

idk????now

thomaster (thomaster):

there's -24 on the left. How do you get it to the right?

OpenStudy (anonymous):

add 24?

thomaster (thomaster):

Yes :) So you'll get 3n=24 now only 1 thing left to do

OpenStudy (anonymous):

3n-24+24=0+24 3n=24 divide 3 answer 8?

thomaster (thomaster):

Very good!

thomaster (thomaster):

Last one: \(\large 5(2-g)=0 \)

OpenStudy (anonymous):

last??? ): ): ): walking me thru them helps alottt for me to understand. don't stop pleaseee)": )';

OpenStudy (anonymous):

distribute 5 10-5g=0 subtract 10

thomaster (thomaster):

See, you already know how to solve it :) What's next? We now have: \(\large -5g=-10\)

OpenStudy (anonymous):

divide -5

OpenStudy (anonymous):

g= -2?

thomaster (thomaster):

when you divide -10 by -5 the -'s cancel eachother out. So the answer is 2 :)

OpenStudy (anonymous):

leaving??? )":

thomaster (thomaster):

Yes, I'd love to help you further but unfortunately I really have to go now :((( Post question 20-26 in a new question and i'm sure someone else will help you with this :P

OpenStudy (anonymous):

4 more??????

OpenStudy (anonymous):

thanks anyways. thanks for medal.. I REALLY appreciate the help, I understand a lot better.. good night